OpenStudy (anonymous):

How do I find the slope of the line which passes through (2, 6) and (0, 2)?

OpenStudy (anonymous):

Try using Geogebra it is very useful!!

OpenStudy (anonymous):

Go to google and type in Geogebra.

OpenStudy (anonymous):

Gradient = (6 - 2) / (2 - 0) = 4 / 2 = 2 Using (0 , 2): y = 2x + 2

OpenStudy (nurali):

\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

http://www.geogebra.org/cms/en/download/ and click what ever operator system your computer has.

OpenStudy (anonymous):

Use the equation y2-y1/x2-x1 So it would be 2-6 / 0-2 And the answer would be (-4,-2)

OpenStudy (anonymous):

But the answer are 2 -2 -3 4

OpenStudy (anonymous):

The answer is 2

OpenStudy (anonymous):

ok thanks